OpenStudy (anonymous):

FAN & MEDAL. - 9th Grade Integrated Mathematics 1. What are the first three terms of the sequence a1 = 1 and an = 4(an-1)^3? A. 1,4, 256 B. 1, 256, 500 C. 4, 32, 108 D. 4, 8, 256

OpenStudy (anonymous):

Here's a picture for better visual on the problem: http://i.gyazo.com/36cbdab5bc67fb4cd3860fcc2d7ee1b0.png

OpenStudy (lyrae):

You substitute an-1 with the value for a1 and evaluate that expression to get a2. You then do the same thing again with a2 to get a3.

OpenStudy (anonymous):

My first guess on the question would be A. 1, 4, 256. I don't really quite understand the equation but I kind of get an idea what you're saying.

OpenStudy (lyrae):

Maby an example will help: \[a_1=3\]\[a_n = 2(a_{n-1})^2\] \[a_2 = 2(a_{2-1})^2 = 2(a_{1})^2 = 2(3)^2 = 2 \times 9 = 18\] \[a_3 = 2(a_{3-1})^2 = 2(a_{2})^2 = 2(18)^2 = 2 \times 324 = 648\]

OpenStudy (lyrae):

And yes it's A :)

OpenStudy (anonymous):

Wow, I'll definitely keep your example in mind. Thank you very much, @Lyrae

OpenStudy (lyrae):

Anytime :)